# Python is Not Always the Best Choice: Embracing Multilingual Program of Thoughts

### v2

@@ Abstract
Program of Thoughts (PoT) is an approach characterized by its executable intermediate steps, which ensure the accuracy of the logical calculations in the reasoning process. Currently, PoT primarily uses Python. However, relying solely on a single language may result in suboptimal solutions and overlook the potential benefits of other programming languages. In this paper, we conduct comprehensive experiments on the programming languages used in PoT and find that no single language consistently delivers optimal performance across all tasks and models. The effectiveness of each language varies depending on the specific scenarios. Inspired by this, we propose a task and model agnostic approach called MultiPoT, which harnesses strength and diversity from various languages. Experimental results reveal that it significantly outperforms Python Self-Consistency. Furthermore, it achieves comparable or superior performance compared to the best monolingual PoT in almost all tasks across all models. In particular, MultiPoT achieves more than 4.6% improvement on average on ChatGPT (gpt-3.5-turbo-0701).
@@
Xianzhen Luo1, Qingfu Zhu1thanks:   Corresponding author, Zhiming Zhang1, Libo Qin3, Xuanyu Zhang2, Qing Yang2, Dongliang Xu2, Wanxiang Che1 1Harbin Institute of Technology, Harbin, China 2Du Xiaoman (Beijing) Science Technology Co., Ltd. 3Central South University, Changsha, China {xzluo, qfzhu, zmzhang, car}@ir.hit.edu.cn {zhangxuanyu, yangqing, xudongliang}@duxiaoman.com {m15840540712}@163.com
@@ 5.1 Comparison among PLs
However, it is important to note that there is no one-size-fits-all language. The gap between PLs is significant when considering each task and model. The performance of each PL is task-dependent. AVG performance does not fully capture the disparity among languages. Java and JavaScript performances of Starcoder differ by only 0.41% on AVG, but by 6.71% on Tabular. While the difference between the best and worst PLs of ChatGPT on AVG is less than 2% in Table 1, there are four tasks whose gap among languages exceeds 6%. Different languages are suitable for different tasks. Table 1 indicates that, except for C++, all PLs excel in at least one task on ChatGPT. Moreover, on ChatGPT, except for JavaScript, each language also ranks as the least effective in at least one task. A language that performs exceptionally well in one task might underperform in another. For instance, R demonstrates superior performance on Date for both Code LLMs and ChatGPT, yet it is the least effective on Appl. and Math.
The performance of each PL is model-dependent. Code LLMs and ChatGPT differ significantly. The results of three Code LLMs are averaged and compared with ChatGPT in Table 1. It shows that C++, JavaScript, and Java, excel on Appl., Math, and Spatial respectively on Code LLMs, but rank second-to-last on ChatGPT. Even within Code LLMs, disparities between models are evident. Figure 3 shows that Code Llama has a clear preference for Java, which keeps the top two ranks across all tasks, yet is not observed on the remaining models. On Deepseek Coder, C++ leads on average, whereas it ranks last on the other models. R ranks second on Spatial on Deepseek Coder, but the worst on the other two Code LLMs.
@@ 5.2 Comparision between Self-Consistency and MultiPoT
Self-Consistency does not eliminate performance disparities between PLs, despite it significantly improving the performance. Table 2 presents the Self-Consistency results. The inherent strength of different languages persist. The optimal PL on each scenario is generally consistent with greedy decoding results, except Python emerges as the superior language on Math on all model. A single language offers limited diversity. When faced with tasks outside its strength, monolingual samples often make the same mistakes repeatedly, resulting in incorrect answers being chosen through voting.
@@ 6.1 Reasoning Ability of Different Languages
In Section 5.1, we note that the ranking of the average performance of PL varies on each model. The language distribution in the pre-training data of Starcoder and Deepseek Coder offers insights into whether data amount impacts reasoning capabilities. Moreover, we are interested in examining whether code generation and reasoning of multilingual ability are aligned. The difference between the two tasks is elucidated in Appendix A.4. To assess code generation ability, we utilize the results of each model on the Multilingual HumanEval benchmark, focusing on the four available languages, excluding R due to a lack of evaluation dataset.
@@ 6.2 MultiPoT Analysis
MutliPoT has the highest coverage rate. Unlike the voting mechanism which requires a majority for the correct answer, the coverage rate only needs the answer to appear in results. Table 3 demonstrates coverage rates on all four models and MultiPoT achieves the highest. The monolingual sampling covers less than the multilingual attempts, highlighting that the strength of different PLs exists. MultiPoT effectively utilizes the strength of different PLs and has the highest upper bound.
@@ A.1 Tasks
Appl. comprises the GSM8K (Cobbe et al., 2021), SVAMP (Patel et al., 2021), and Asdiv (Miao et al., 2020) datasets. These datasets contain elementary-level math problems set in specific application scenarios, focusing on mathematical abstraction and modeling skills, with relatively low difficulty. Since they are the same type of questions, we merge them in one task. Math, consisting of the transformed MATH (Hendrycks et al., 2021) dataset, whose the answers to the problems are expressed using LaTeX. It’s too hard to construct prompts in other languages that meet all the requirements, we select those that can be calculated to a single number, excluding problems with interval or formula-based answers. The filtered results is shown in Table 7.
@@ A.3 Error Analyse
A further categorisation of the types of RE is conducted. We classify all REs into eight error types. Redeclaration represents duplicate naming of variables. Division by Zero represents the denominator in the division is zero. Illegal Output represents the answer can not be parsed or converted correctly. Time Limit Error represents the program runs out of time and sometimes it is due to stack space overflow. Compile Error often means there are some syntax error in the program. Undefined Identifier includes Undefined Variables and Undefined Functions, which means the variables or functions are not defined before they are used. Variable Type Error indicates that the types of variables are mismatched when they are involved in some operations, for example addition or division. Table 15 shows the proportion of different RE types for Deepseek Coder across five tasks and five languages. Table 16 presents the proportion of various RE types for four LLMs on Appl. across all languages. Deepseek Coder and the Appl. task are selected because the languages have the most similar performance on them. The results demonstrate that even in scenarios where languages exhibit similar performance, the proportions of RE differ significantly among languages. For instance, the RE rate on ChatGPT’s Appl. of R and C++ differs by only 0.02%, yet Illegal Output account for 82.46% of C++ errors, in comparison to only 24.71% for R. Given that each prompt is accurate, the differing error distributions are attributable to the intrinsic characteristics of the languages, thereby demonstrating their diversity and the non-repetitive nature of their errors.

### v4

@@ Abstract
Program of Thoughts (PoT) is an approach characterized by its executable intermediate steps, which ensure the accuracy of the logical calculations in the reasoning process. Currently, PoT primarily uses Python. However, relying solely on a single language may result in suboptimal solutions and overlook the potential benefits of other programming languages. In this paper, we conduct comprehensive experiments on the programming languages used in PoT and find that no single language consistently delivers optimal performance across all tasks and models. The effectiveness of each language varies depending on the specific scenarios. Inspired by this, we propose a task and model agnostic approach called MultiPoT, which harnesses strength and diversity from various languages. Experimental results reveal that it significantly outperforms Python Self-Consistency. Furthermore, it achieves comparable or superior performance compared to the best monolingual PoT in almost all tasks across all models. In particular, MultiPoT achieves more than 4.6% improvement on average on ChatGPT (gpt-3.5-turbo-0701) [^1] Code and data are released at https://github.com/Luowaterbi/MultiPoT.
@@
Xianzhen Luo1, Qingfu Zhu1thanks:   Corresponding author, Zhiming Zhang1, Libo Qin2, Xuanyu Zhang3, Qing Yang3, Dongliang Xu3, Wanxiang Che1 1Harbin Institute of Technology, Harbin, China 2Central South University, Changsha, China 3Du Xiaoman (Beijing) Science Technology Co., Ltd. {xzluo, qfzhu, zmzhang, car}@ir.hit.edu.cn lbqin@csu.edu.cn {zhangxuanyu, yangqing, xudongliang}@duxiaoman.com
@@ 4.2 Tasks
4.3 Metric
Remaining consistent with previous work Chen et al. (2022); Gao et al. (2023), the metric is accuracy. For tasks whose ground truth are real numbers (Appl./Math), the answer is considered correct if its difference from the ground truth is less than 1e-3; for tasks with string-type ground truth (Date/Tabular/Spatial), the answer is considered correct only if it is exactly the same as the ground truth.
@@ 5.1 Comparison among PLs
However, it is important to note that there is no one-size-fits-all language. The gap between PLs is significant when considering each task and model.
The performance of each PL is task-dependent. AVG performance does not fully capture the disparity among languages. Java and JavaScript performances of Starcoder differ by only 0.41% on AVG, but by 6.71% on Tabular. While the difference between the best and worst PLs of ChatGPT on AVG is less than 2% in Table 1, there are four tasks whose gap among languages exceeds 6%. Different languages are suitable for different tasks. Table 1 indicates that, except for C++, all PLs excel in at least one task on ChatGPT. Moreover, on ChatGPT, except for JavaScript, each language also ranks as the least effective in at least one task. A language that performs exceptionally well in one task might underperform in another. For instance, R demonstrates superior performance on Date for both Code LLMs and ChatGPT, yet it is the least effective on Appl. and Math.
The performance of each PL is model-dependent. Code LLMs and ChatGPT differ significantly. The results of three Code LLMs are averaged and compared with ChatGPT in Table 1. It shows that, on Appl., C++ performs best on Code LLMs but ranks second-to-last on ChatGPT; on Math, JavaScript excels on Code LLMs but similarly ranks second-to-last on ChatGPT; and on Spatial, Java ranks second-highest on Code LLMs (with only a 0.05% less than C++) but is second-to-last on ChatGPT. Even within Code LLMs, disparities between models are evident. Figure 3 shows that Code Llama has a clear preference for Java, which keeps the top two ranks across all tasks, yet is not observed on the remaining models. On Deepseek Coder, C++ leads on average, whereas it ranks last on the other models. R ranks second on Spatial on Deepseek Coder, but the worst on the other two Code LLMs.
@@ 5.2 Comparision between Self-Consistency and MultiPoT
Self-Consistency does not eliminate performance disparities between PLs, despite it significantly improving the performance. Table 2 presents the Self-Consistency results. The inherent strength of different languages persist. The optimal PL on each scenario is generally consistent with greedy decoding results, except Python emerges as the superior language on Math on all model. The weaknesses of each language is further amplified. For example, on Date of Deepseek Coder, C++ already had the lowest performance in greedy decoding, and Self-Consistency increases this gap even more. As a result, C++ shifts from the highest average performance in greedy decoding on Deepseek Coder to the lowest in Self-Consistency, despite remaining the best on Appl., Tabular, and Spatial. A single language offers limited diversity. When faced with tasks outside its strength, monolingual samples often make the same mistakes repeatedly, resulting in incorrect answers being chosen through voting.
@@ 6.1 Reasoning Ability of Different Languages
In Section 5.1, we note that the ranking of the average performance of PL varies on each model. The language distribution in the pre-training data of Starcoder and Deepseek Coder offers insights into whether data amount, defined as the percentage of each language in the pre-training corpus, impacts reasoning capabilities. Moreover, we are interested in examining whether code generation and reasoning of multilingual ability are aligned. The difference between the two tasks is elucidated in Appendix A.4. To assess code generation ability, we utilize the results of each model on the Multilingual HumanEval benchmark, focusing on the four available languages, excluding R due to a lack of evaluation dataset.
@@
#Shots 0 3
RE WA AC RE WA AC
Python 17.33 29.97 52.71 2.72 32.14 65.14
R 55.81 15.45 28.74 2.08 34.47 63.44
C++ 59.61 14.90 25.48 0.34 30.87 68.79
Java 14.86 33.25 51.89 0.54 32.07 67.38
JavaScript 69.26 13.77 16.96 1.74 32.41 65.84
Table 3: Comparison for different PLs under zero-shot (0) and 3-shot (3) demonstrations. RE represents Runtime Error. WA means Wrong Answer. AC represents ACcuracy.
@@
Zero-shot reasoning ability shows considerable inconsistency when compared to 3-shot reasoning ability. Table 3 presents the results of zero-shot and 3-shot experiments using Code Llama 34B on Appl., where AC denotes accuracy, and incorrect outcomes are further classified into Runtime Errors (RE) and Wrong Answers (WA). The results reveal particularly steep declines in R, C++, and JavaScript, largely driven by a significant increase in RE. This suggests that different PLs exhibit varying levels of sensitivity to shot settings. Two prominent error patterns emerge from the zero-shot outputs: (1) LLM frequently generates repetitive comments until reaching the maximum sequence length, and (2) LLM generates CoT without corresponding executable code. These observations highlight the importance of high-quality, language-specific demonstrations; only with effective demonstrations can the model fully harness the reasoning capabilities of different PLs.
@@ 6.2 MultiPoT Analysis
MutliPoT has the highest coverage rate. Unlike the voting mechanism which requires a majority for the correct answer, the coverage rate is the percentage of questions that have at least one correct answer in five candidate answers in the dataset. For example, if the candidate answers to a question are “(6 5 5 5 7)” and the ground truth is “7”, the question is covered. Coverage rate can be considered as an upper bound because this metric represents the proportion of all potentially solvable problems, assuming there exists a selection mechanism better than the current voting mechanism. Table 4 demonstrates coverage rates on all four models and MultiPoT achieves the highest. The monolingual sampling covers less than the multilingual attempts, highlighting that the strength of different PLs exists. MultiPoT effectively utilizes the strength of different PLs and has the highest upper bound.
@@
Date Table Spatial
CoT 82.38 91.28 97.70
PoT 79.40 97.28 97.95
MultiPoT 80.22 98.66 99.10
Table 8: Comparison between Self-Consistency of CoT, PoT (Python), and MultiPoT. CoT results are based on Deepseek LLM v2, while PoT and MultiPoT are based on Deepseek Coder v2.
MultiPoT is better than CoT Self-Consistency. To compare the performance of CoT and PoT in scenarios where precise mathematical calculations are not required, we conduct experiments on Date, Table, and Spatial—using Deepseek LLM v2 (a 405B MoE LLM) for CoT and Deepseek Coder v2 (continually pretrained from Deepseek LLM v2) for PoT. The results, shown in Table 8, indicate that PoT achieves better Self-Consistency than CoT on Table and Spatial, with MultiPoT further improving performance. On Table, the improvement demonstrates the advantage of PoT in understanding and reasoning over structured data. On Date, however, PoT slightly underperforms CoT, which is primarily due to PoT interpreting the difference between two dates as exclusive, while natural language typically considers it inclusive. Nevertheless, the results suggest that PoT remains valuable in scenarios where precise calculations are unnecessary, and MultiPoT continues to be effective.
@@ 7 Conclusion
Acknowledge
We gratefully acknowledge the support of the National Natural Science Foundation of China (NSFC) via grant 62236004, 62206078, 62441603 and 62476073 and the support of Du Xiaoman (Beijing) Science Technology Co., Ltd.
@@ A.1 Tasks
Appl. comprises the GSM8K (Cobbe et al., 2021), SVAMP (Patel et al., 2021), and Asdiv (Miao et al., 2020) datasets. These datasets contain elementary-level math problems set in specific application scenarios, focusing on mathematical abstraction and modeling skills, with relatively low difficulty. Since they are the same type of questions, we merge them into one task. Math, consisting of the transformed MATH (Hendrycks et al., 2021) dataset, whose answers to the problems are expressed using LaTeX. It’s too hard to construct prompts in other languages that meet all the requirements, we select those that can be calculated to a single number, excluding problems with interval or formula-based answers. The filtered results are shown in Table 9.
@@ A.3 Error Analysis
A further categorization of the types of RE is conducted. We classify all REs into eight error types. Redeclaration represents duplicate naming of variables. Division by Zero represents the denominator in the division is zero. Illegal Output represents the answer can not be parsed or converted correctly. Time Limit Error represents the program runs out of time and sometimes it is due to stack space overflow. Compile Error often means there are some syntax error in the program. Undefined Identifier includes Undefined Variables and Undefined Functions, which means the variables or functions are not defined before they are used. Variable Type Error indicates that the types of variables are mismatched when they are involved in some operations, for example addition or division. Table 17 shows the proportion of different RE types for Deepseek Coder across five tasks and five languages. Table 18 presents the proportion of various RE types for four LLMs on Appl. across all languages. Deepseek Coder and the Appl. task are selected because the languages have the most similar performance on them. The results demonstrate that even in scenarios where languages exhibit similar performance, the proportions of RE differ significantly among languages. For instance, the RE rate on ChatGPT’s Appl. of R and C++ differs by only 0.02%, yet Illegal Output account for 82.46% of C++ errors, in comparison to only 24.71% for R. Given that each prompt is accurate, the differing error distributions are attributable to the intrinsic characteristics of the languages, thereby demonstrating their diversity and the non-repetitive nature of their errors.
Upon further analysis of generated contents, several common failure patterns emerge:
  • Date Calculation: PoTs often misinterpret the difference between two dates as exclusive, contrary to natural language conventions, where the interval is typically inclusive. Nevertheless, R demonstrates comparable performance to CoT on Date, indicating its potential for temporal reasoning tasks.
  • Output Content: PoTs frequently respond to yes/no questions by outputting attributes instead of directly answering the question. For example, when asked, ‘On the desk, there is a teal pen and a yellow textbook. Is the textbook yellow?’, the correct answer is ‘yes’, but PoT might respond with ‘yellow’.
  • Demonstration Constraint: Demonstrations are less restrictive for JavaScript, as it tends to output extra information beyond what is required, including descriptive sentences and variables, even when only the final answer is needed.
  • Syntax Preference: C++ and Java tend to leverage language-specific constructs like for-loops, often reaching correct solutions in pattern recognition problems. In contrast, other languages that attempt step-by-step calculations may add or omit steps, leading to errors.
  • Resource Constraints: Low-resource languages like R may cause the model to call non-existent packages, particularly in complex mathematical problems that rely on external tools.